# Graph-SCP: Accelerating Set Cover Problems with Graph Neural Networks

### v2

@@ Abstract
Machine learning (ML) approaches are increasingly being used to accelerate combinatorial optimization (CO) problems. We investigate the Set Cover Problem (SCP) and propose Graph-SCP, a graph neural network method that augments existing optimization solvers by learning to identify a much smaller sub-problem that contains the solution space. Graph-SCP uses both supervised learning from prior solved instances and unsupervised learning aimed at minimizing the SCP objective. We evaluate the performance of Graph-SCP on synthetically weighted and unweighted SCP instances with diverse problem characteristics and complexities, and on instances from the OR Library, a canonical benchmark for SCP. We show that Graph-SCP reduces the problem size by 60–80% and achieves runtime speedups of up to 10x on average when compared to Gurobi (a state-of-the-art commercial solver), while maintaining solution quality. This is in contrast to fast greedy solutions that significantly compromise solution quality to achieve guaranteed polynomial runtime. We showcase Graph-SCP’s ability to generalize to larger problem sizes, training on SCP instances with up to 3,000 subsets and testing on SCP instances with up to 10,000 subsets.
Introduction
@@
A growing area of research uses machine learning (ML) solutions to combinatorial optimization (CO) problems. These ML solution can be divided into two categories: (1) learning an end-to-end model that generates a feasible solution (e.g., Dai et al. (2017)) and (2) learning a non-end-to-end model that predicts branching heuristics and aids an existing CO solver (e.g., Gasse et al. (2019)). Bengio, Lodi, and Prouvost (2021) provide a comprehensive survey of methods used in each category. Here, we focus on the latter approach of using ML to aid conventional CO solvers.
@@
Our focus is on the set cover problem (SCP), which is an NP-hard CO problem. In SCP, one is given the “universe,” which is a set of elements $\{1,2,\dots m\}$. One is also given a collection of $n$ sets whose union corresponds to the universe. To solve the problem, one must identify the smallest sub-collection of the $n$ sets whose union is equal to the universe.
@@
We use a graph neural network (GNN) model to accelerate the runtime of SCP solutions without sacrificing solution quality. Concretely, we propose Graph-SCP, where we cast an instance of SCP as a graph and learn a GNN to predict a subgraph that encapsulates the solution space. The nodes of this subgraph are passed into a conventional CO solver such as Gurobi (Gurobi Optimization, LLC 2023), which is a state-of-the-art commercial solver. The smaller size of the input problem to traditional solvers enables them to run faster. Graph-SCP achieves between 60–80% reduction in input problem size, which leads to runtime improvements of up to $10$x on average, all the while maintaining solution quality across a range of SCP instance characteristics.
@@
Traditional off-the-shelf solvers swiftly generate incumbent solutions, but may require additional time to reach optimality. In our comparative analysis, we examine the incumbent solutions generated by both Graph-SCP and Gurobi, revealing that the search space reduction component of Graph-SCP allows a traditional solver to find incumbent and finally, optimal solutions faster. Our contributions are as follows:
@@
  • We provide insights into how features from alternative representations of an SCP instance can be used to improve Graph-SCP’s search. Our findings show that features derived from a hypergraph representation of an SCP instance leads to the largest speedup.
  • We present a comprehensive evaluation of Graph-SCP, encompassing SCP instances ranging across various densities, sizes, and costs (weighted vs. unweighted), as well as instances from the OR Library (Beasley 1990). We demonstrate Graph-SCP’s ability to generalize across various problem characteristics. We compare against two benchmark state-of-the-art solvers: Gurobi (Gurobi Optimization, LLC 2023) and SCIP (Bestuzheva et al. 2023); and discuss how various modeling choices affect the performance of Graph-SCP.
Background
Figure 1(A) shows an example of a SCP instance. We adopt an alternate formulation for SCP. Given a binary matrix $A\in\mathbf{R}^{m\times n}$, SCP is defined as covering all $m$ rows (denoting the elements of the universe) by the minimum cost subset of the $n$ columns (denoting the collection of sets whose union corresponds to the universe). Costs of each column are represented in the column vector $c\in\mathbf{R}^{n\times 1}$. The goal is to find the assignment vector $x\in\mathbf{R}^{n\times 1}$, where
@@
$\min_{x}\sum_{{j\in n}}c_{j}x_{j}\textrm{~{}~{}s.t. }\ \sum_{j\in n}A_{ij}x_{j}\geq 1,\text{ }\forall i\in m,x_{j}\in\{0,1\}.$ (2)
Given such a matrix, we can define its density as $d=\frac{q}{m\times n}$, where $q$ is the number of non-zero entries in the matrix $A$ (Lan, DePuy, and Whitehouse 2007).
Graph Abstractions for SCP
By treating the covering matrix as an adjacency matrix with elements in the universe as rows and sets as columns, we can represent the SCP instance as a directed graph. We chose a directed tripartite graph (see Figure 1(B)). A node representing the universe is connected to each element node, and the element nodes are connected to the sets in which they appear.
We also represent the SCP instance as a hypergraph. Here, each element of the universe is a vertex and each set is a hyperedge. (See Figure 1(C); the hyperedges are demarcated by bounding boxes around the associated element nodes.) This hypergraph representation offers an additional lens through which to explore the connections between elements and sets. For example, it enables an investigation into the structural characteristics of an SCP instance by examining features inherent to the hypergraph structure (such as hyperedge dependent vertex weights; see Figure 1(D)).
Figure 1: (A) A simple example of a set cover problem with equal cost subsets. (B) The set cover instance represented using a bipartite graph abstraction. This graph is passed into a GNN which then predicts whether nodes (i.e. subsets) belong to the SCP solution (green highlighted nodes) or not. The prediction values for Layer 0 and Layer 1 nodes (white-colored nodes) are discarded, since only Layer 2 nodes contribute to the solution. (C) The hypergraph representation for the same set cover instance. Each subset corresponds to a hyperedge shown as bounding boxes around the element nodes. (D) We define hyperedge dependent weights for each element of the universe as the inverse of the size of its associated hyperedge as shown.
Graph Neural Networks (GNNs).
GNNs combine node wise updates with message updates from neighboring nodes. Throughout this work, we utilize GraphSAGE (Hamilton, Ying, and Leskovec 2017), a message passing GNN. We discuss other popular GNN architectures in Table 4 of the Appendix. Formally, given a graph $G=(V,E)$ with $V$ nodes, $E$ edges, and $\vec{x_{v}}$ node features, GraphSAGE computes node representations as:
@@
$h_{v}^{k}\leftarrow\sigma(W^{k}\cdot CONCAT(h_{k}^{k-1},h^{k}_{N(v)}),$ (5)
where $k$ is the number of message passing layers, $h^{k}_{v}$ is the hidden representation of node $v$ at layer $k$, $W^{k}$ is the weight matrix at layer $k$, $N(v)$ is the set of node $v$’s neighbors, and $AGGREGATE$ can be any permutation invariant aggregation method like sum or average.
Proposed Method: Graph-SCP
Figure 2: Graph-SCP takes as input a SCP instance represented as a graph. The graph is augmented with a set of features on its nodes ($G_{f}$). A GNN is trained on this graph to predict a subgraph of nodes that likely contains the solution to the SCP instance. At runtime, the GNN is used (only once) to generate predictions. Graph-SCP picks nodes at a predefined percentile threshold as the subgraph. If the objective criteria are not met, the threshold is reduced, thus selecting a larger subgraph.
Figure 2 provides an overview of Graph-SCP, which takes as input a SCP instance represented as a graph $G$ (a.k.a. a SCP graph). Graph-SCP augments $G$ with a set of features on its nodes (described in the next section). Then, a GNN is trained on the augmented graph, $G_{f}$, to predict whether a subgraph (i.e., a set of nodes) is likely to appear in the solution. The output of the GNN is subsequently given to a thresholding procedure that selects which subgraphs should be given to the solver. If the solver cannot find a solution, Graph-SCP reduces the threshold for selecting subgraphs. This process continues until a solution is found.
Augmenting a SCP Graph with Features
@@
  • Cost: For the “Universe” node in layer 0 of the graph (see Figure 1B), the cost is set to 0. For each of the “Elements” nodes in layer 1 of the graph, the cost is also set to 0. For each of the “Columns” nodes in layer 2 (representing a set of nodes), this feature is set to the cost of the corresponding set (which is part of the SCP input).
  • Cover: This feature is the cardinality of the set of elements represented by the node. For each of the “Columns” nodes in layer 2, it is the size of the set. For the “Elements” nodes in layer 1 of the graph, it is set to 1. For the “Universe” node, it is set to 0. This feature is similar to the one used by the standard greedy approximation algorithm for SCP (Chvatal 1979).
  • Random Walks with Restarts (RWR) (Tong, Faloutsos, and Pan 2006): This feature represents the relative importance of nodes based on the structure of the graph. Our RWR algorithm starts a random walk from the “Universe” node in layer 0 and restarts the walk (with probability $p=0.45$) from the same node (see Figure 1(B)). (The restart probability is a hyperparameter. We set the value to 0.45 since the SCP graph is shallow.) When the walker reaches stationary distribution, the score between each node in layers 1 and 2 and the universe node is used as a node feature. This feature value for the “Universe” node is set as 0.
  • Degree-based features: For each node in the SCP graph (Figure 1(B)), we capture its degree and the average degree of its neighbors.
  • Hypergraph features: Figure 1(C) shows an SCP instance cast as a hypergraph. Given a hypergraph $H(V,E,\omega,\gamma)$, where each vertex $v\in V$ has hyperedge-dependent weights $\gamma_{e}(v)$ and hyperedge weights $\omega(e)$ (denoting the cost of each set), we compute its Laplacian matrix (Chitra and Raphael 2019). The algebraic connectivity of $H$ (i.e., the second smallest eigenvalue of the Laplacian matrix of $H$) is denoted by $\mu(H)$. The contribution of an individual hyperedge $e$ to the algebraic connectivity of $H$ can be quantified as $c(e)=\mu(H)-\mu(H-e)\geq 0,\forall e\in E.$
    According to the definition of algebraic connectivity, $H$ is connected when $\mu(H)>0$. In the context of SCP, connectedness implies the coverage of every element in the universe. For an unweighted SCP instance modeled as a hypergraph, the optimal solution corresponds to a connected hypergraph with the fewest hyperedges. This optimization objective can be expressed as the goal to maximize algebraic connectivity, while minimizing the number of hyperedges. A greedy approximation that iteratively removes hyperedges with the smallest $c(e)$ will return an approximate solution to an SCP instance.
    Compared to a standard greedy algorithm for SCP, which relies only on the hyperedge degree as a heuristic (i.e., number of uncovered elements in a set), features derived from the hypergraph Laplacian capture richer information than the hyperedge degree alone (see Figure 3).
@@
    To generate hypergraph features for the “Elements” and “Columns” nodes, we use the following procedure. First, each “Elements” node is assigned a weight equal to the inverse of its hyperedge degree – i.e., the weight of an element within a subset corresponds to the inverse of the set size (see Figure 1(D)). Second, we define two probability transition matrices in order to generate the Laplacian matrix of the SCP hypergraph with edge-dependent vertex weights (Chitra and Raphael 2019). The first probability transition matrix is for transitions from one “Elements” node to another via a hyperedge (a “Columns” node); and the second matrix is for transitions from one hyperedge to another via an “Elements” node. Third, we compute the eigenvalues of the two Laplacian matrices. These eigenvalues become the hypergraph-based features of the “Elements” and “Columns” nodes, respectively.
Using a GNN to Predict Membership in the Optimal Solution
@@
Chen et al. (2022) proved that given a dataset of SCP instances, a GNN can approximate the optimal solution of the linear program relaxation of the SCP instance. Building on this observation, we formalize reducing the SCP problem size as a subgraph selection learning task. In particular, our Graph-SCP uses a message passing GNN to learn the following binary classification task: predict which “Columns” nodes (see Figure 1(B)) are likely to be part of the optimal solution. Formally, given a graph $G=(V,E)$ with node features $X$, we learn a GNN function $f(G,X)$ to predict a binary classification vector $y\in\mathbb{R}^{|V|}$, where 1 indicates the nodes in the graph that are likely to contain the SCP solution and 0 otherwise.

### v3

@@ Abstract
Machine learning (ML) approaches are increasingly being used to accelerate combinatorial optimization (CO) problems. We investigate the Set Cover Problem (SCP) and propose Graph-SCP, a graph neural network method that augments existing optimization solvers by learning to identify a smaller sub-problem that contains the solution space. Graph-SCP uses both supervised learning from prior solved instances and unsupervised learning to minimize the SCP objective. We evaluate the performance of Graph-SCP on synthetically weighted and unweighted SCP instances with diverse problem characteristics and complexities, and on instances from the OR Library, a canonical benchmark for SCP. We show that Graph-SCP reduces the problem size by 60–80% and achieves runtime speedups of up to 10x on average when compared to Gurobi (a state-of-the-art commercial solver), while maintaining solution quality. This is in contrast to fast greedy solutions that significantly compromise solution quality to achieve guaranteed polynomial runtime. We showcase Graph-SCP’s ability to generalize to larger problem sizes, training on SCP instances with up to 3,000 subsets and testing on SCP instances with up to 10,000 subsets.
1 Introduction
@@
A growing area of research explores machine learning (ML) solutions to combinatorial optimization (CO) problems. These ML solutions can be divided into two categories: (1) learning end-to-end models that generates feasible solutions (e.g., [11]) and (2) learning non-end-to-end models that predict branching heuristics and aids existing CO solvers (e.g., [14, 31]). Bengio et al. [3] provide a comprehensive survey of methods used in each category. Here, we focus on the latter approach and use ML to aid conventional CO solvers.
@@
We focus on the set cover problem (SCP), an NP-hard CO problem. In SCP, one is given the “universe,” which is a set of elements $\{1,2,\dots m\}$. One is also given a collection of $n$ sets whose union corresponds to the universe. To solve the problem, one must identify the smallest sub-collection of the $n$ sets whose union is equal to the universe.
@@
We use a graph neural network (GNN) model to accelerate the runtime of SCP solvers without sacrificing solution quality. Concretely, we propose Graph-SCP, where we cast an instance of SCP as a graph and learn a GNN to predict a subgraph that encapsulates the solution space. The nodes of this subgraph are passed into a conventional CO solver such as Gurobi [16] (a state-of-the-art commercial solver). This method of reducing the size of the input problem and passing it to traditional solvers enables them to run faster. Graph-SCP achieves between 60–80% reduction in input problem size, which leads to runtime improvements of up to 1010x on average, while maintaining solution quality across a range of SCP instance characteristics.
@@
Traditional solvers swiftly generate incumbent solutions, but require additional time to reach optimality. In our comparative analysis, we examine the incumbent solutions generated by both Graph-SCP and Gurobi, revealing that the reduction of the search space by Graph-SCP allows the solver to find incumbent and finally, optimal solutions faster. Our contributions are as follows:
@@
  • We provide insights into how features from alternative representations, such as a hypergraph representation of an SCP instance can be used to improve Graph-SCP’s search. Our findings show that features derived from this hypergraph representation of the SCP instance lead to the largest speedup.
  • We present a comprehensive evaluation of Graph-SCP, with SCP instances ranging across various densities, sizes, and costs (weighted vs. unweighted), as well as instances from the OR Library [2]. We demonstrate Graph-SCP’s ability to generalize across problem characteristics. We compare against two benchmark state-of-the-art solvers, Gurobi [16] and SCIP [4], and discuss how various modeling choices affect the performance of Graph-SCP.
2 Background
Figure 1(A) shows an example of an SCP instance. We adopt an alternate formulation for SCP. Given a binary matrix $A\in\mathbb{R}^{m\times n}$, SCP is defined as covering all $m$ rows (denoting the elements of the universe) by the minimum cost subset of the $n$ columns (denoting the collection of sets whose union corresponds to the universe). Costs of each column are represented in the column vector $c\in\mathbb{R}^{n}$. The goal is to find the assignment vector $x\in\mathbb{R}^{n}$, where
@@ 2 Background
$\min_{x}\sum_{{j\in n}}c_{j}x_{j}\textrm{~~s.t. }\ \sum_{j\in n}A_{ij}x_{j}\geq 1,\text{ }\forall 1\leq i\leq m,x_{j}\in\{0,1\}.$ (2)
Given such a matrix, we can define its density as $d=\frac{q}{m\times n}$, where $q$ is the number of non-zero entries in the matrix $A$ [24].
Graph Abstractions for SCP.
By treating the covering matrix as an adjacency matrix, with elements in the universe as rows and sets as columns, we can represent the SCP instance as a directed bipartite graph (see Figure 1(B)). A node representing the universe is connected to each element node, and the element nodes are connected to the sets in which they appear.
We also represent the SCP instance as a hypergraph. Here, each element of the universe is a vertex and each set is a hyperedge (Figure 1(C)). The hyperedges are demarcated by bounding boxes around the associated element nodes. This hypergraph representation offers an additional lens through which to explore the connections between elements and sets. For example, it enables an investigation into the structural characteristics of an SCP instance by examining features inherent to the hypergraph structure (such as hyperedge dependent vertex weights in Figure 1(D)).
Figure 1: (A) A simple example of a set cover problem with equal cost subsets. (B) The set cover instance represented using a bipartite graph abstraction. This graph is passed into a GNN which then predicts whether nodes (i.e. subsets) belong to the SCP solution (green highlighted nodes). The prediction values for Layer 0 and Layer 1 nodes (white-colored nodes) are discarded, since only Layer 2 nodes contribute to the solution. (C) The hypergraph representation for the same set cover instance. Each subset corresponds to a hyperedge shown as bounding boxes around the element nodes. (D) For each hyperedge $e$, we define hyperedge dependent weights $\gamma_{e}(v)$ for each element $v$ of the universe as the inverse of the size of its associated hyperedge as shown.
Graph Neural Networks.
GNNs combine node-wise updates with message updates from neighboring nodes. Throughout this work, we utilize GraphSAGE [17], a message passing neural network. Formally, given a graph $G=(V,E,X)$ with nodes $V$, edges $E$, and node features $X$, GraphSAGE computes node representations as:
@@
[…]
where $\vec{x_{v}}$ is the node feature vector for node $v$, $k$ is the number of message passing layers, $h^{k}_{v}$ is the hidden representation of node $v$ at layer $k$, $W^{k}$ is the weight matrix at layer $k$, $N(v)$ is the set of node $v$’s neighbors, and $AGGREGATE$ can be any permutation invariant aggregation method like sum or average.
3 Proposed Method: Graph-SCP
Figure 2: Graph-SCP takes an SCP instance represented as a graph as input. The graph is augmented with a set of features on its nodes ($G_{f}$). A GNN is trained on this graph to predict a subgraph of nodes that likely contains the solution to the SCP instance. At runtime, the GNN is used (only once) to generate predictions. Graph-SCP picks nodes at a predefined percentile threshold as the subgraph. If the objective criteria are not met, the threshold is reduced, thus selecting a larger subgraph.
Figure 2 provides an overview of Graph-SCP, which takes as input an SCP instance represented as a graph $G$ (a.k.a. an SCP graph). Graph-SCP augments $G$ with a set of features on its nodes (described in the next section). Then, a GNN is trained on the augmented graph, $G_{f}$, to predict whether a subgraph (i.e., a set of nodes) is likely to appear in the solution. The output of the GNN is subsequently given to a thresholding procedure that selects which subgraphs should be given to the solver. If the solver cannot find a solution, Graph-SCP reduces the threshold and selects a larger subgraph. This process continues until a solution is found.
3.1 Augmenting a SCP Graph with Features
@@
Cost:
We assign costs of 0 to the “Universe” node in layer 0 and “Element” nodes in layer 1 of the graph (Figure 1(B)). The “Columns” nodes in layer 2 (representing subsets) have their costs set according to the input problem.
Cover:
This feature is the cardinality of the set of elements represented by the node. For each of the “Columns” nodes in layer 2, it is the size of the set. For the “Elements” nodes in layer 1 of the graph, it is set to 1. For the “Universe” node, it is set to 0. This feature is similar to the one used by the standard greedy approximation algorithm for SCP [10].
Random Walks with Restarts (RWR) [34]:
This feature represents the relative importance of nodes based on the structure of the graph. Our RWR algorithm starts a random walk from the “Universe” node (layer 0) and restarts the walk from the same node (see Figure 1(B)). When the walker reaches stationary distribution, the score between each node in layers 1 and 2 and the universe node is used as a node feature. This feature value for the “Universe” node is set as 0. The restart probability is a hyperparameter and is set to 0.45 since the SCP graph is shallow.
Degree-based features:
For each node in the SCP graph (Figure 1(B)), we capture its degree and the average degree of its neighbors.
Hypergraph features:
Figure 1(C) shows an SCP instance cast as a hypergraph. For a hypergraph $H(V,E,\omega,\gamma)$, each hyperedge $e$ has weights $\omega(e)$ (denoting the cost of each set). The elements within each hyperedge $e$, denoted by $v\in V$ have hyperedge-dependent weights $\gamma_{e}(v)$. These hyperedge-dependent weights correspond to the inverse of the hyperedge degree as shown in Figure 1(D). Given this hypergraph, we compute its Laplacian matrix [9]. The algebraic connectivity of $H$, i.e., the second smallest eigenvalue of the Laplacian matrix of $H$, is denoted by $\mu(H)$. The contribution of an individual hyperedge $e$ to the algebraic connectivity of $H$ can be quantified as $c(e)=\mu(H)-\mu(H-e)\geq 0,\forall e\in E.$ According to the definition of algebraic connectivity, $H$ is connected when $\mu(H)>0$. In the context of SCP, connectedness implies the coverage of every element in the universe. For unweighted SCP instances modeled as hypergraphs, the optimal solution corresponds to a connected hypergraph with the fewest hyperedges. The optimization objective can be expressed as maximizing algebraic connectivity and minimizing the number of hyperedges. A greedy approach that iteratively removes hyperedges with the smallest $c(e)$ will return an approximate solution to the SCP instance.
Compared to a standard greedy algorithm for SCP which relies only on the hyperedge degree as a heuristic (i.e., number of uncovered elements in a set), features derived from the hypergraph Laplacian capture richer information than the hyperedge degree alone (see Figure 3).
@@
To generate hypergraph features for the “Elements” and “Columns” nodes, we use the following procedure. First, each “Elements” node is assigned a weight equal to the inverse of its hyperedge degree – i.e., the weight of an element within a subset corresponds to the inverse of the set size (see Figure 1(D)). Second, we define two probability transition matrices in order to generate the Laplacian matrix of the SCP hypergraph with edge-dependent vertex weights [9]: (i) a matrix for transitions from one “Elements” node to another via a hyperedge (a “Columns” node), and (ii) a matrix for transitions from one hyperedge to another via an “Elements” node. Third, we compute the eigenvalues of the two Laplacian matrices. These eigenvalues become the hypergraph-based features of the “Elements” and “Columns” nodes, respectively.
3.2 Using a GNN to Predict Membership in the Optimal Solution
@@
Chen et al. [8] proved that given a dataset of SCP instances, a GNN can approximate the optimal solution of the linear program relaxation of the SCP instance. Building on this observation, we formalize reducing the SCP problem size as a subgraph selection learning task. In particular, our Graph-SCP uses a message passing GNN to learn the following binary classification task: predict which “Columns” nodes (see Figure 1(B)) are likely to be part of the optimal solution. Formally, given a graph $G=(V,E)$ with node features $X$, we learn a GNN function $f(G,X)$ to predict a binary classification vector $y\in\mathbb{R}^{|V|}$, where 1 indicates the nodes in the graph that are likely to contain the SCP solution and 0 otherwise.
Objective Function.
Graph-SCP’s objective function consists of supervised and unsupervised components. The supervised component is a standard binary cross entropy loss that uses labels generated by Gurobi, whereas the unsupervised component directly minimizes the objective value of the linear program relaxation of the SCP instance. Formally, let $A\in\mathbb{R}^{m\times n}$ be the binary covering matrix and $c\in\mathbb{R}^{n\times 1}$ be the cost vector. Then the loss function is given by
[…]
Inference.
The trained model outputs continuous values between 0 and 1 for each “Column” node from layer 2. To select a set of nodes, Graph-SCP selects nodes at a predefined percentile threshold and passes the selected nodes into a CO solver. Our experiments use Gurobi and SCIP; however, other solvers of choice can be used. Subsequently, Graph-SCP checks if the objective value objobj returned by the solver with the reduced problem size is at least a user-defined value $u_{obj}$, i.e., $obj\leq u_{obj}$. If not, the percentile threshold is reduced (by a fixed decrement size, set to 10 in our experiments), thus selecting a larger set of nodes to pass into the solver. Graph-SCP uses solutions from each iteration to warm-start the solver for the next iteration. This iterative refinement process mirrors Column Generation [15] methods used in tackling large-scale CO problems.
In our experiments, we set $u_{obj}$ to be the optimal objective value, thereby forcing Graph-SCP to run until the optimal value is reached. It is important to note that $u_{obj}$ serves as a bound rather than a specific solution, indicating the proximity to optimality required for the solution; an aspect often specified in practical applications. In real-world scenarios where a predefined limit is not available, an additional iteration can be run to determine if further improvements are possible. Due to the warm-starting mechanism, where each iteration builds on the results of the previous one, these additional runs incur relatively low computational cost. Furthermore, we demonstrate that in situations constrained by runtime rather than a specific objective limit, Graph-SCP consistently identifies superior incumbent solutions compared to Gurobi (see Figure 6).
Node Feature Importance.
Figure 3 demonstrates the value of the node features described in the previous section for Graph-SCP’s prediction task. For each “Elements” node and each “Columns” node, we extract the node embeddings from the penultimate layer of their GNN. To visualize these embeddings, we use PaCMAP [38] and map the embeddings in 2 dimensions. To quantify how well various sets of features can separate solution nodes from non-solution nodes, we measure the average distance of a solution node to other solution nodes (intra distance) and the average distance of a solution node to non-solution nodes (inter distance) in the GNN embedding space. Ideally, the intra distance is small, the inter distance is large and difference between inter and intra distance is large. With this definition in place, we see in Figure 3 that separability is the highest when using all of the aforementioned node features, i.e., subplots (A) and (E) where the difference between inter and intra distances is the largest.
Figure 3: Visualization of node embeddings in two dimensions using PaCMAP [38]. The embeddings are extracted from the penultimate layer of Graph-SCP’s message passing GNN with different sets of node features. The features used are displayed above each column. Each row shows embeddings for a different type of SCP instance (see Table 1). To quantify how well various sets of features can separate solution nodes from non-solution nodes, we measure the average distance of a solution node to other solution nodes (intra distance) and the average distance of a solution node to non-solution nodes (inter distance) in the GNN embedding space. Observe how (A) and (E) have the largest inter distance with small intra distances, which indicates that Cost, Cover, RWR, degree-based, and hypergraph-based features are useful in separating solution nodes from non-solution nodes.
4 Experimental Setup
GNN Architecture.